madison999:

Nick bought a music player. The price was $172, and the sales tax rate was 7 percent. How much sales tax did Nick pay when he bought the music player? A. $7.20 B. $12.04 C. $12.45

xXAikoXx:

The sales tax can be found by multiplying Base price × $172 $172 × 0.07 = $12.04 So that means, he paid $12.04 tax for the item.

madison999:

@xxaikoxx wrote:
The sales tax can be found by multiplying Base price × $172 $172 × 0.07 = $12.04 So that means, he paid $12.04 tax for the item.
thank you

xXAikoXx:

@madison999 wrote:
@xxaikoxx wrote:
The sales tax can be found by multiplying Base price × $172 $172 × 0.07 = $12.04 So that means, he paid $12.04 tax for the item.
thank you
My pleasure.